# Unusual switch from low-temperature T-quadratic resistivity in the underdoped pseudogap phase of cuprate superconductors to low-temperature T-linear resistivity in the overdoped strange-metal phase

### v1

@@ Abstract
The transport experiments demonstrate unambiguously a dramatic switch from the low-temperature linear in temperature (T-linear) resistivity in the overdoped strange-metal phase to the low-temperature quadratic in temperature (T-quadratic) resistivity in the underdoped pseudogap phase of cuprate superconductors, however, a consensus on the origin of this unusual switch is still lacking. Here the low-temperature resistivity in the underdoped pseudogap phase of cuprate superconductors is investigated using the Boltzmann transport equation. The low-temperature resistivity originates from the electron umklapp scattering mediated by the spin excitation. However, the dominant contribution to the low-temperature resistivity mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of a momentum dependent spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic in the low temperatures below $T_{\rm scale}$ with the strength of the T-quadratic resistivity that weakens as the doping is raised. However, in the overdoped regime, the resistivity is T-linear in the low temperatures above $T_{\rm scale}$. The result in this paper together with the recent study on the electrical transport in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime resulting from the opening of a momentum dependent spin pseudogap. Concomitantly, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
@@ I Introduction
It has become very clear that cuprate superconductors are among the most complicated systems studied in condensed matter physics Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . The complications stems mainly from the fact that the parent compound of cuprate superconductors is a form of non-conductor called a Mott insulator Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . Superconductivity then is achieved by chemically introducing charge carriers to this Mott insulator, which leads to that the physical properties of cuprate superconductors significantly depend on the extent of doping, and the regimes have been classified into the underdoped, optimally doped, and overdoped, respectively. After intensive investigations over more than three decades, a substantial amount of reliable and reproducible data has been accumulated by using many probes Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 , which reveals that the most of the nonconventional and unique features of cuprate superconductors are observed in the normal-state above the superconducting (SC) transition temperature $T_{\rm c}$, where the deepest mysteries lie. In particular, below a pseudogap crossover temperature $T^{*}$, which can be well above $T_{\rm c}$ in the underdoped regime, the physical response can be well interpreted in terms of the formation of a pseudogap by which it means a large suppression of the electronic density of states at the electron Fermi surface (EFS) Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . This is why in the underdoped regime, the phase above $T_{\rm c}$ but below $T^{*}$ is so-called as the pseudogap phase. The pseudogap in the underdoped regime was first discovered through Knight-shift of the magnetic susceptibility experiments Warren89 ; Alloul89 , where the pseudogap is some kind of the spin pseudogap, meaning a strong reduction in the magnetic susceptibility through the particle-hole correlation. Subsequently, the presence of the pseudogap was confirmed by a series of experimental measurements taken with a wide variety of techniques Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . Moreover, these experiments also show clearly that the pseudogap exists in both the spin and charge channels Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . For example, angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling spectroscopy measure the charge channel Loeser96 ; Norman98 , while nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) detect the spin channelFujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 . On the other hand, in the overdoped regime, the normal-state is characterized by a number of the anomalous low-temperature properties Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 in the sense that they do not fit in with the conventional Fermi-liquid theory Schrieffer64 ; Abrikosov88 ; Mahan81 . This has led to the normal-state in the overdoped regime being refereed to as the strange-metal phaseKeimer15 .
@@
Among the notable characteristics in the normal-state of cuprate superconductors, the most iconic feature has to be the low-temperature resistivity Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 . In the overdoped regime, a series of transport measurements revealed a linear in temperature (T-linear) resistivity at low temperaturesKeimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , which is in a striking contrast to the conventional Fermi-liquid behaviour found in the conventional superconductorsSchrieffer64 ; Abrikosov88 ; Mahan81 . However, in the underdoped regime, the early transport measurements indicated that the opening of the pseudogap in the spin excitation spectrum leads to a striking deviation from the low-temperature T-linear behaviour of the resistivityGurvitch87 . In particular, the subsequent transport experiments observed a reduced resistivity below $T^{*}$, near the temperature where the spin pseudogap, as seen by NMR Knight-shift and spin-relaxation rate, openedBucher93 ; Ito93 ; Nakano94 ; Ando01 . Thus the low-temperature resistivity anomaly coincides roughly with $T^{*}$ and these experimental results suggest that the decrease in the low-temperature below $T^{*}$ is caused by the reduced electron scattering by the spin excitations resulting from the opening the spin pseudogapBucher93 ; Ito93 ; Nakano94 ; Ando01 . Lately, the low-temperature resistivity was confirmed experimentally in a quadratic temperature dependenceAndo04a ; Ando04 ; Lee05 ; Proust08 , as would be expected from the conventional Fermi-liquid theory, for a wide doping range in the underdoped regime. More specifically, the recent transport experiments demonstrate unambiguously a clear and dramatic switch from the low-temperature T-linear resistivity in the overdoped regime to the purely quadratic in temperature (T-quadratic) resistivity in the underdoped regime Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 , where the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) decreases with the increase of doping. In spite of the considerable variation in crystal structures and impurity-scattering effects among different families of cuprate superconductors, these experimental results Gurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 ; Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 also show clearly that as the low-temperature T-linear resistivity in the overdoped regimeLegros19 ; Ayres21 ; Grisso21 , the low-temperature T-quadratic resistivity also is a universal feature in the underdoped regime. In this case, some crucial questions are raised: (i) why the low-temperature resistivity exhibits a dramatic switch from the T-linear resistivity in the overdoped regime to the T-quadratic resistivity in the underdoped regime? (ii) whether the pseudogap is correlated with this dramatic switch or not? (iii) is there a common electron scattering mediated by the same bosonic mode that is responsible for both the low-temperature T-quadratic resistivity in the underdoped regime and low-temperature T-linear resistivity in the overdoped regime?
@@
The theoretical explanation for this unusual switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime along with the underlying scattering mechanism is clearly a big challenge. For a understanding of the nature of the low-temperature T-linear resistivity in the overdoped regime, several scenarios have been proposed Varma89 ; Varma16 ; Varma20 ; Damle97 ; Sachdev11 ; Zaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 ; Lee21 . However, these scenarios are just as diverse as the mechanism of superconductivity. In the marginal Fermi-liquid phenomenologyVarma89 ; Varma16 ; Varma20 , a single T-linear scattering rate is posited to account for the T-linear resistivity. In particular, the T-linear resistivity was interpreted as a consequence of the scale invariant physics near to the quantum critical point (QCP) Varma20 ; Damle97 ; Sachdev11 . With the close relation to the physics of QCP, the T-linear resistivity has been interpreted in terms of the Planckian dissipationZaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 . Moreover, it has been shown recently that the low-temperature resistivity originates from the umklapp scattering between electrons by the exchange of a critical boson propagatorLee21 , where the umklapp scattering rate is T-linear in low temperatures near the umklapp point, which induces a low-temperature T-linear resistivity. These studiesLee21 together with many othersHonerkamp01 ; Hartnoll12 thus suggest that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity. On the other hand, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regime Puchkov96 , then the spin pseudogap seen in NMR below $T^{*}$ would naturally account for the deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . Moreover, the mechanism for the resistivity in the underdoped regime has been proposedHussey03 ; Rice17 , where the electron umklapp scattering above $T^{*}$ leads to the T-linear resistivity Rice17 . As the temperatures fall below $T^{*}$, the pseudogap opens, and then restricts the available umklapp scattering channels, leading to a crossover between the T-linear resistivity at the temperature above $T^{*}$ and the traditional T-quadratic form of the scattering from the Fermi pockets at the temperature below $T^{*}$. However, up to now, the microscopic origin of the low-temperature T-quadratic resistivity in the underdoped regime and its connection with the pseudogap are still much debated, and no explanation of the unusual switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime has been made so far starting from a microscopic theory.
@@
In the very recent studyMa23 , we have investigated the nature of the low-temperature T-linear resistivity in the overdoped strange-metal phase. In our scenario, the scattering rate is arisen from the umklapp scattering between electrons by the exchange of the effective spin propagator, where the dominant contribution to the resistivity mainly comes from the antinodal umklapp scattering. This umklapp scattering rate scales linearly with temperature in the low temperatures, which then naturally generates a low-temperature T-linear resistivity. In this paper, we study the low-temperature electrical transport in the underdoped pseudogap phase along with this line, where in a striking difference to the case in the overdoped regime, the spin excitation energy dispersion is anisotropically renormalized due to the opening of a momentum dependent spin pseudogap. In particular, the density of the spin excitations at around the antinodal is heavily reduced by the antinodal spin pseudogap. In this case, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this low-temperature scale $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ is dropped down with the enhancement of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the T-quadratic resistivity strength weakens as the doping is raised. However, in the overdoped regime, the resistivity is instead T-linear in the low temperatures above $T_{\rm scale}$. The present results combined with the recent findingsMa23 on the electrical transport in the overdoped regime therefore reveal that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependent spin pseudogap lowers the density of the spin excitations at around the antinodal region in the response to the umklapp scattering, which reduces the strength of the intense umklapp scattering from the electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. Consequently, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
@@
This paper is organized as follows. In the next section, we first provide a brief summary for the derivation of the full spin and electron propagators, where (i) the antinodal spin pseudogap decreases with the increase of doping in the underdoped regime, and then it abruptly vanishes at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. Following this full spin propagator and the reconstructed EFS, the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived using the Boltzmann transport equation. The discussions of the quantitative characteristics of the low-temperature resistivity in the underdoped pseudogap phase are given in Section III, where we show that the scattering rate in the whole doping regime is predominantly governed by the antinodal umklapp scattering. In Section IV, we close the presentation with a summary and discussion of the main results.

### v2

# Low-temperature T2 resistivity in the underdoped pseudogap phase versus T-linear resistivity in the overdoped strange-metal phase of cuprate superconductors

@@ Abstract
The transport experiments demonstrate a dramatic switch from the low-temperature linear in temperature (T-linear) resistivity in the overdoped strange-metal phase of cuprate superconductors to the low-temperature quadratic in temperature (T-quadratic) resistivity in the underdoped pseudogap phase, however, a consensus on the origin of this unusual switch is still lacking. Here the resistivity in the underdoped pseudogap phase of cuprate superconductors is investigated using the Boltzmann transport equation. The resistivity originates from the electron umklapp scattering mediated by the spin excitation, however, the dominant contribution mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of a momentum dependent spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Moreover, this $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the strength of the T-quadratic resistivity weakens as the doping is raised. However, in the overdoped regime, the resistivity is T-linear at the low temperatures above $T_{\rm scale}$. The results in this paper together with the recent study on the resistivity in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime resulting from the opening of a momentum dependent spin pseudogap.
@@ I Introduction
It has become very clear that cuprate superconductors are among the most complicated systems studied in condensed matter physics Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . The complications stem mainly from a fact that the parent compound of cuprate superconductors is an antiferromagnetic (AF) insulator Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . Superconductivity then is achieved by chemically introducing charge carriers to this AF insulator, which leads to that the physical properties mainly depend on the extent of doping, and the regimes have been classified into the underdoped, optimally doped, and overdoped, respectively. After intensive studies about four decades, a substantial amount of reliable and reproducible data has been accumulated by using many probes Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 , which reveals that the most of the unusual features of cuprate superconductors are observed in the normal-state. In particular, below a pseudogap crossover temperature $T^{*}$, which can be well above the superconducting (SC) transition temperature $T_{\rm c}$ in the underdoped regime, the physical response can be well interpreted in terms of the formation of a pseudogap by which it means a large suppression of the electronic density of states on the electron Fermi surface (EFS) Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . This is why in the underdoped regime, the phase above $T_{\rm c}$ but below $T^{*}$ is so-called as the pseudogap phase. The pseudogap in the underdoped regime was first discovered through Knight-shift of the magnetic susceptibility experiments Warren89 ; Alloul89 , where the pseudogap is some kind of the spin pseudogap, meaning a strong reduction in the magnetic susceptibility through the particle-hole correlation. Subsequently, the presence of the pseudogap was confirmed by a series of experimental measurements taken with a wide variety of techniques Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . Moreover, these experiments also show clearly that the pseudogap exists in both the spin and charge channels Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . For example, angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling spectroscopy measure the charge channel Loeser96 ; Norman98 , while nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) detect the spin channelFujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 . On the other hand, in the overdoped regime, the normal-state is characterized by a number of the anomalous low-temperature properties Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 in the sense that they do not fit in with the conventional Fermi-liquid theory (FL) Schrieffer64 ; Abrikosov88 ; Mahan81 . This has led to the normal-state in the overdoped regime being refereed to as the strange-metal phaseKeimer15 .
@@
Among the notable characteristics in the normal-state, the most iconic feature has to be the electrical transportHussey08 ; Kastner98 ; Timusk99 ; Hufner08 . In the overdoped regime, a series of transport measurements revealed a T-linear resistivityKeimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , where (i) the linear temperature term often extends to low temperatures of a few kelvin; and (ii) the linear temperature term persists to high temperatures with the same slope all the way to the lowest temperature. In this paper, we shall not address the case at high temperature, but instead focus on the low-temperature T-linear resistivity, which is in a striking contrast to the conventional FL behaviour in the conventional superconductorsSchrieffer64 ; Abrikosov88 ; Mahan81 . However, in the underdoped regime, the early transport measurements indicated that the opening of the pseudogap in the spin excitation spectrum leads to an obvious deviation from the low-temperature T-linear behaviour of the resistivityGurvitch87 . In particular, the subsequent transport experiments observed a reduced resistivity below $T^{*}$, near the temperature where the spin pseudogap, as seen by NMR Knight-shift and spin-relaxation rate, openedBucher93 ; Ito93 ; Nakano94 ; Ando01 . Thus the low-temperature resistivity anomaly coincides roughly with $T^{*}$ and these experimental results suggest that the decrease in the low-temperature resistivity below $T^{*}$ is caused by the reduced electron scattering by the spin excitations resulting from the opening the spin pseudogapBucher93 ; Ito93 ; Nakano94 ; Ando01 . Lately, the low-temperature resistivity was confirmed experimentally in a quadratic temperature dependenceAndo04a ; Ando04 ; Lee05 ; Proust08 , as would be expected from the conventional FL theorySchrieffer64 ; Abrikosov88 ; Mahan81 , for a wide doping range in the underdoped regime. More specifically, the recent transport experiments demonstrate a clear and dramatic switch from the low-temperature T-linear resistivity in the overdoped regime to the purely T-quadratic resistivity in the underdoped regimeCooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , where the strength of the T-quadratic resistivity decreases with the increase of doping. In spite of the considerable variation in crystal structures and impurity-scattering effects among different families of cuprate superconductors, these experimental results Gurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 ; Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 also show that as the low-temperature T-linear resistivity in the overdoped regime Legros19 ; Ayres21 ; Grisso21 , the T-quadratic resistivity also is a universal feature in the underdoped pseudogap phase. In this case, some crucial questions are raised: (i) why the low-temperature resistivity exhibits the T-quadratic behaviour in the underdoped regime, with an exotic crossover to the T-linear behaviour in the overdoped regime? (ii) whether the pseudogap is correlated with this exotic crossover or not? (iii) is there a common electron scattering mediated by the same bosonic mode that is responsible for both the low-temperature T-quadratic resistivity in the underdoped regime and low-temperature T-linear resistivity in the overdoped regime?
@@
The theoretical explanation for this exotic crossover along with the underlying scattering mechanism is clearly a big challenge. For a understanding of the nature of the low-temperature T-linear resistivity in the overdoped regime, several scenarios have been proposed Varma89 ; Varma16 ; Varma20 ; Damle97 ; Sachdev11 ; Zaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 ; Lee21 . However, these scenarios are just as diverse as the mechanism of superconductivity. In the marginal FL phenomenologyVarma89 ; Varma16 ; Varma20 , a single T-linear scattering rate is posited to account for the T-linear resistivity. In particular, the T-linear resistivity was interpreted as a consequence of the scale invariant physics near to the quantum critical point (QCP) Varma20 ; Damle97 ; Sachdev11 . With the close relation to the physics of QCP, the T-linear resistivity has been interpreted in terms of the Planckian dissipationZaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 . Moreover, it has been shown recently that the low-temperature T-linear resistivity originates from the umklapp scattering between electrons by the exchange of a critical boson propagator Lee21 . These studiesLee21 together with many othersHonerkamp01 ; Hartnoll12 thus suggest that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity. On the other hand, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regime Puchkov96 , then the spin pseudogap seen in NMR below $T^{*}$ would naturally account for a deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . In particular, the mechanism for the resistivity in the underdoped regime has been proposedHussey03 ; Rice17 , where the electron umklapp scattering above $T^{*}$ leads to the T-linear resistivityRice17 . As the temperatures fall below $T^{*}$, the pseudogap opens, and then restricts the available umklapp scattering channels, leading to a crossover between the T-linear resistivity at the temperature above $T^{*}$ and T-quadratic form of the scattering from the Fermi pockets at the temperature below $T^{*}$. Moreover, the low-temperature resistivity has been studied based on the two-particle self-consistent approachBergeron11 , where it has been shown that the resistivity is linear at low temperatures in the overdoped regime, however, in the underdoped regime, the resistivity may display a T-quadratic behaviour at temperatures below $T^{*}$. However, up to now, the T-quadratic resistivity in the underdoped pseudogap phase and its connection with the pseudogap is still the subject of much study, and the microscopic origin of the clear switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime is hotly debated.
@@
In the very recent studyMa23 , we have studied the nature of the low-temperature T-linear resistivity in the overdoped regime. In our scenario, the scattering rate arises from the umklapp scattering between electrons by the exchange of the effective spin propagator, where the dominant contribution mainly comes from the antinodal umklapp scattering. This umklapp scattering rate scales linearly with temperature in the low temperatures, which then naturally generates a low-temperature T-linear resistivity. In this paper, we study the low-temperature resistivity in the underdoped pseudogap phase along with this line, where in a striking difference to the case in the overdoped regime, the spin excitation energy dispersion is anisotropically renormalized due to the opening of a momentum dependent spin pseudogap. In particular, the density of the spin excitations at around the antinodal is heavily reduced by the antinodal spin pseudogap. In this case, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this low-temperature scale $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ is dropped down with the enhancement of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the T-quadratic resistivity strength weakens as the doping is raised. However, in the overdoped regime, the resistivity is instead T-linear in the low temperatures above $T_{\rm scale}$. The present results combined with the recent resultsMa23 on the low-temperature T-linear resistivity in the overdoped regime therefore reveal that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependent spin pseudogap lowers the density of the spin excitations at around the antinodal region, which reduces the strength of the intense umklapp scattering from the electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate.
@@
This paper is organized as follows. In the next section, we first show that (i) the antinodal spin pseudogap decreases with the increase of doping in the underdoped regime, and then it abruptly vanishes at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. Following this reconstructed EFS, the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived using the Boltzmann transport equation. The discussions of the quantitative characteristics of the low-temperature resistivity in the underdoped pseudogap phase are given in Section III, where we show that the scattering rate in the whole doping regime is predominantly governed by the antinodal umklapp scattering. In Section IV, we close the presentation with a summary and discussion of the main results. In Appendix A, we present the details for the derivation of the full spin and full electron propagators.